Bandar Industries Berhad of Malaysia manufactures sporting equipment. One of the company’s products, a football helmet for the North American market, requires a special plastic. During the quarter ending June 30, the company manufactured 3,700 helmets, using 2,220 kilograms of plastic. The plastic cost the company $16,872.

According to the standard cost card, each helmet should require 0.51 kilograms of plastic, at a cost of $8.00 per kilogram.

Bandar Industries Berhad of Malaysia manufactures
Answer #1

number of helmet 3700
standrd kilogram of plastic per helmet 0.51
total standard kilogram allowed 1887
standard cost per kg $8
total standard cost $15,096
actual cost incurred $16,872
total material variance $1,776


MATERIAL PRICE VARIANCE =(ACTUAL QUANTITY *ACTUAL PRICE)-(ACTUAL QUANTITY*STANDARD PRICE) =(2220*7.6)-(2220*8)= $-888


MATERIAL QUANTITY VARIANCE =(AQ*SP)-(SQ*SP)= (2220*8)-(1887*8)=$ 2664